Science Report – November 17th

Science Report
Camila Castillo
17 Novermber 2017

The research project about the effect of Streptomyces spp. strain could not be finished properly. The first problem was related with the lack of medium needed for finishing this project. Nutrient broth and Nutrient agar were the ones I worked with, so there are better results with these. I decided to buy these elements via Internet, but I had to wait a week. After that week, there was little time to start the project. I decided to start working right away, but sadly, the medium ended up contaminated. I will need to check if the transportation of the strain was done properly, though. It would seem that that was the main problem. Maybe it couldn’t resist as much as I expected. Microbiology projects tend to have these kind of problems, though. As a beginner in the microbiological world, maybe I still need to learn some more about microbiology techniques to work an entire project in just two weeks! I will be taking with me some soil samples to Peru. I hope that the enviroment provided by my lab at Lima will help me to finish this project.
Despite the kind of failure of my main project, I managed to learn some things about martian agriculture. First of all, the soil isn’t as simple as I expected. It hardens and it prevents the water to filter through it. It has little to non-organic matter. Somehow I thought that my mere strain could help with these problems, but I realized that this is a problem I will need to adress if I want to finish this (which I will). I tried to search for a technique that didn’t involved mixing two types of soil, but both didn’t work. It seems that in the end I will have to do the mixing.
However, I also managed to do some projects besides my main one. I tried to see the effects of the dishwater on the seeds we had, and I managed to do a nice diagram about some samples and their growth in a couple of days. I also decided to describe the microbes in different plant soils we managed to collect on the EVAs. I’ve been amazed by the shape and forms I managed to get on my plates. I will try to take some of these samples to Lima and process them, to know exactly which microbes are in there.

EVA Report – November 17th

EVA #10 Report

Today’s EVA began at 10:00 AM, involving Camila, Julio, and I. The Survivalman crew was filming our EVA, including the planning process, communications tests, putting on the spacesuits, and our decompression. The goal of our EVA was to go to Lith Canyon, collect three regolith samples, and return to the Hab. Each of the EVA crew boarded an ATV, with Les Stroud on his own ATV, and his camera crew on Deimos. I had Camila lead the way out so we could keep our ATV caravan nice and tight. The trip there took about an hour, because they needed footage of us along several different stretches of Cow Dung Rd. We would hold our position as they moved ahead, then we would move forward once they signaled that they were ready.

Once we arrived at the end of the Lith Canyon road, we proceeded on foot for about a kilometer westward. We had to be careful with our route finding as this was an area that none of us had been to before. We made sure to photograph the landscape at certain points so it would be easier to find our location later. Once we descended into the canyon, we made our way to an exposure with red and white regolith layers, taking one sample from each type. Once finished we walked further down the canyon where there were several eolian rock formations. There we collected our last regolith sample from a white exposure, turned around, and headed back the way we came. We spent about 2 hours walking in the Lith Canyon area, returning to the ATVs by around 1:20 PM. They didn’t need any more footage for the trip back, so we made our way back to the Hab by around 1:50 PM.

Today’s EVA was abnormally warm, especially all the walking we had to do. This was the longest EVA that I’ve had to do on foot, and was very challenging. I’d estimate that we walked a total of about 2 km over uneven terrain. It’s amazing how heavy and uncomfortable the suits felt on the first EVA when we just walked around the Hab. Now I feel a lot better in the EVA suit, even when covering much longer distances. Knowing this, I’m glad we didn’t plan any really long EVAs during our first few days of Sim, as it does take some time to get used to those suits. I’d definitely recommend that crew members take their time getting used to the suits, as their bulkiness can be surprising.

For our EVA, we managed to obtain the regolith samples we needed and the Survivorman crew was able to record our EVA operations.

This is all I have to report for today’s EVA.

Have a great night Mission Support!

 

Prepared by Brandon Ferguson

EVA Officer

EVA Report – November 16th

EVA #9 Report
Today’s EVA would’ve started at 11 AM and ended at 2 PM. Our goal would’ve been to go to Lith Canyon, collect samples from the Salt Wash Member, and return to the Hab. Our EVA for today did not take place, since we needed additional time to clean up the Hab before Survivorman arrives. Also we needed to figure out our transfer arrangements so that we’ll be able to get to Salt Lake City in time to catch their flight. EVA request coming soon!
Prepared by Brandon Ferguson
EVA Officer

GreenHab Report – November 16th

Sol 13 – 16/11/201

Green Hab Officer Atila Meszaros

Ambient: heating 10°C

Ambient with window/door open: door opened at 10:20; closed at 17:34

Shade cloth on/off: on

Average temperature:

Time 10:36 17:06
Temperature inside (°C) 27 15
Temperature outside (°C) 10 8
% Humidity 16 16

Changes to crops:  Crops keep growing!

Daily water usage for crops: 2 galons

Time(s) of watering for crops:  10:36 and 17:06

Changes to research plants: controls keep growing! No signs of germination on the analog soil. The emergency day turned the analog soil like clay, again.

Daily watering and amount of water used: 1 ½ L approximately

Narrative:

Hi, Sylvain! Atila, here. I hope you are having an amazing night.

Kevin’s crops are growing pretty solid, Julio already sent a picture. It’s crazy! I wanted to ask what’s the plan for them.

I thought that the seeds on the Petri dishes were ready to be planted on the analog soil, again. They just needed another one more day. The quinoa seeds are not germinating on the control, either. I’m guessing that the day I didn’t water them was enough for them from drying. That almost killed all the project! I will need to continue with my experiment on my university with the analog soil to get solid results.

I think that’s pretty much all, if you need more information about the GH I will be glad to send it to you.

Thank you for being there, Sylvain!

Have a beautiful night on Earth.

Support/supplies needed: none

Operations Report – November 16th

Crew 182 Operations Report  16Nov2017

SOL: 13

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER: Finally, we found the problem. The new water heater is not installed, yet.

CHARGERS OF THE SUITS: Two of the chargers are working, but we moved them and now it is hard for us to connect them, again.

 

Generator—  It worked the whole night since 16:30pm of 15Nov2017 until 12:00 pm of 16Nov2017.  Turn on at 16:50 pm 16Nov2017.  Operand.

Solar—

READING FROM 16 NOV 2017

 

TIME SOC READINGS
12:00 pm 100%
13:00 pm 98%
14:00 pm 91%
15:00 pm 84%
16:20 pm 80%
16:50 pm 75%

 

Diesel – 50 %

Propane – 49 psi

Gasoline (5 Gallon containers for ATV) –  2 ½ gallons.

Water (trailer) –  0 gallons

Water (static) – 300  gallons

Trailer to Static Pump used –   No

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  60 gallons

Toilet tank emptied:  yes

ATV’s Used: No

Oil Checked:  No

Atv # Fuel Used Gals:  0  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 0 hours

Deimos used:  No

Hours: 0

Beginning charge:  –

Ending charge:  –

Currently charging: Yes
HabCar used? No

Notes and Comments:

I would call today, the day of cleaning. Everyone here is tired. Literally, we cleaned every hole here. There was too much dust. Today in the morning, the commander went to search transportation to return to the Earth.  We are sad to leave Mars, and at the same time, happy for returning at home.

From another hand, survivor man is coming. We are so excited¡

 

Summary of Hab Operations:
We cleaned the whole hab. We found, a couple of glasses and gloves, behind the area of the tools.

Summary of GreenHab operations:

We cleaned the greenhouse. We found too much dust. The two wheelbarrows are in the greenhouse, now.

Summary of ScienceDome operations:

The science dome is finally cleaned. In fact, we found tons of dust under the carpet.

We are using the lamps of the Hab, for the laboratory.

Summary of health and safety issues:

Now, we can say the station is in equilibrium with our health and safety.

Questions, concerns and requests to Mission Support:

We need gasoline for the ATV’S, please. URGENT

Sol Summary – November 16th

MDRS Daily Summary Report for Sol 13.

Summary Title: Organizing MDRS.

Mission Status: Preparing to leave MDRS. We did some improvements in MDRS fixing some issues in the physical facilities.

Sol Activity Summary: We just concluded the cleaning procedures of the entire installation, also cleaning the GreenHab and the laboratory. We developed many different activities organizing and fixing the facilities of MDRS.

Look Ahead Plan: Tomorrow, probably will happen an EVA, but We are just waiting the SurvivorMan to decide about would be interesting to him.

Anomalies in work: Nope.

Weather: 17o C, 62o F.

Crew Physical Status: Fine.

EVA: We planed an EVA today, but not was possible. We planed to go to Lith Canyon, collect samples from the Salt Wash Member, and return to the Hab. Our EVA for today did not take place, since we needed additional time to clean up the Hab before Survivorman arrives.

 

Reports to be file:

Commanders Report

Daily Summary Report

Operation Report

GreenHab Report

HSO Report

XO Report

Crew-photos Crew 182 16Nov2017.

 

Support Requested: Nope.

Sol Summary – November 15th

MDRS Daily Summary Report for Sol 12.

Summary Title: Rescuing the ATV.

Mission Status: We developed many kinds of activities today: preparing meals, doing research, practicing exercises and rescuing the ATV.

Sol Activity Summary: Great part of our day was consumed with the rescue of the ATV in the in southeast of Skyline Ridge area. They depart at 12h20m and arrived at 15h30m. Julio and Camila worked in a proposal of guidelines to the Sustainable Development Goals (SDG) during their times aboard of the MDRS. All the crew members developed their specific research.

Look Ahead Plan: Tomorrow, will happen the EVA #9. Starting  11h and finishing 14h. Will participate in the EVA: Brandon (EVA planning, Navigation, Sample Collection), Julio (rock collection and photography) and Carmen (Sample Collection).

Anomalies in work: During the morning part of the crew moves to Skyline Ridge area to rescue the red ATV. Julio broke the simulation between 16h30m-17h30m (1 hour) for 1) to do photos of the full MDRS station from a little distance showing all of the buildings and 2) walking exercise as part of his research related with health and personal sustainability. Julio communicated the external activity to Brandon (EVA Lead).

Weather: 8o C, 46o F.

Crew Physical Status: Fine.

EVA: No EVA today.

 

Reports to be file:

Commanders Report

Daily Summary Report

Scientific Report

Operation Report

GreenHab Report

HSO Report

XSO Report

Crew-photos Crew 182 15Nov2017.

 

Support Requested: Nope.

Operations Report – November 15th

Crew 182 Operations Report  15Nov2017

SOL: 12

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER: Finally, we found the problem. The new water heater is not installed yet.

CHARGERS OF THE SUITS: Two of the chargers are working, but we moved them and now it is hard for us to connect them, again.

 

Generator—  It worked the whole night since 17:05pm of 14Nov2017 until 8:00 am of 15Nov2017.  Turn on at 16:30 pm 15Nov2017.  Operand.

Solar—

READING FROM 15 NOV 2017

 

TIME SOC READINGS
9:30 am 100%
10:30 am 100%
11:30 am 100%
12:30 pm 100%
13:30 pm 100%
14:35 pm 98%
15:30 pm 90/%
16:00 pm 85%
16:30 pm 80%
17:30 pm 100%

 

Diesel – 50 %

Propane – 49 psi

Gasoline (5 Gallon containers for ATV) –  2 ½ gallons.

Water (trailer) –  0 gallons

Water (static) – 350  gallons

Trailer to Static Pump used –   No

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  62 gallons

Toilet tank emptied:  yes

ATV’s Used: Yes

Oil Checked:  No

Atv # Fuel Used Gals:  1  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 2 ½  hours

Deimos used:  Si

Hours: 2 ½

Beginning charge:  – 14:20

Ending charge:  –

Currently charging: Yes
HabCar used? No

Notes and Comments:

We love the new ATV’s and Rovers. Thanks¡

Summary of Hab Operations:
In the morning, when we waked up, we received four more  ATV’s. It is nice, of all them work. Also today, we went to recover the ATV left yesterday in the desert. Also, we needed to charge it, because it does not have battery. We connected it with another ATV, and it works.

Summary of GreenHab operations:

We are cleaning and organizing this area. It is a disaster, but we are working on it.

Summary of ScienceDome operations:

 

We are using the lamps of the Hab, for the laboratory.

Summary of health and safety issues:

Atila is better now.  All of us are better that yesterday.

Questions, concerns and requests to Mission Support:

We need gasoline for the ATV’S, please. URGENT